OpenStudy (indiiavery):

Check my Answers, Please. I'll fan and Medal! 1. What is the slope of the line represented by the equation? y=x/6-5? A. 1/6 B. 6 <-- My Answer C. 5.6 D. –5 2. What is the slope of the line represented by the equation 6x –3y = 4 ? A.2 B.1/2 C.-1/2 <-- My answer D.-2 3. What is the simplified form of the expression? 15y – 3(4y + 10) A. 3y – 30 B.3y + 10 C.27y – 30 D.27y + 10 <-- My answer 4. What is the slope of the line that contains (9,5) and (1,10) My answer : -5,8

OpenStudy (anonymous):

k12?

OpenStudy (indiiavery):

Yes.

OpenStudy (elonasushchik):

All correct! Way too go.

OpenStudy (indiiavery):

Thank you...

jimthompson5910 (jim_thompson5910):

1 through 3 are incorrect

jimthompson5910 (jim_thompson5910):

if you meant to say -5/8 for #4, then you are correct

OpenStudy (elonasushchik):

I was thinking that too but wasn't sure ..

OpenStudy (indiiavery):

Thank you @jim_thompson5910

OpenStudy (anonymous):

y = mx + b if m is the slope and the equation is the same as y=(1/6)x - 5 the answer is A not B. 2: 6x-3y=4 you have to isolate the y to find this -3y=-6x+4 y=2x-(4/3) the slope is 2 3: 15y-3(4y+10) you distribute the -3 first 15y - 12y - 30 then combine like terms 3y-30 4: for this problem you can use the formula y_2 minus y_1 divided by x_2 minus x_1 dont forget: (x,y) (10-5)/(1-9) so the slope is 5/(-8)

OpenStudy (indiiavery):

Thank you @Hally85 !

OpenStudy (anonymous):

No prob, hope that helped :)